Determine if the following sets are linearly dependent, or independent.



(i) {1,sin(x),cos(x)}



(ii) {sin2(x),cos(2x),cos2(x)}


1
Expert's answer
2022-01-04T11:25:12-0500

Let us determine if the following sets are linearly dependent, or independent.


(i) "\{1,\sin(x),\cos(x)\}"


Since the Wronskian


"W(1,\sin(x),\cos(x))=\begin{vmatrix}
1 & \sin(x) & \cos(x)\\
0 & \cos(x) & -\sin(x)\\
0 & -\sin(x) & -\cos(x)
\end{vmatrix}
=
\begin{vmatrix}

 \cos(x) & -\sin(x)\\
 -\sin(x) & -\cos(x)
\end{vmatrix}
\\=-\cos^2(x)-\sin^2(x)=-1\ne 0,"


we conclude that the set "\{1,\sin(x),\cos(x)\}" is linearly independent.


(ii) "\{\sin^2(x),\cos(2x),\cos^2(x)\}"


Taking into account that "\cos(2x)=1\cdot\cos^2(x)-1\cdot\sin^2(x)," we conclude that this set is linearly dependent.